OpenStudy (anonymous):

find the product for the following and please show all work (1/7x^2 - 2/5)(1/2x + 1/2

OpenStudy (ajprincess):

(1/7x^2 - 2/5)(1/2x + 1/2) =1/2x(1/7x^2 - 2/5)+ 1/2(1/7x^2 - 2/5) =1/2x*1/7x^2-1/2x*2/5+1/2*1/7x^2-1/2* 2/5 See if u can continue frm this.

OpenStudy (campbell_st):

just use a grid